# I-FailSense: Towards General Robotic Failure Detection with Vision-Language Models

### v2

@@ I Introduction
Empirically, we demonstrate that I-FailSense outperforms zero-shot SOTA VLMs in detecting semantic misalignment errors reaching an accuracy of 90% in our proposed datasets built from simulated trajectories. Remarkably, I-FailSense generalizes beyond its training data and achieves strong performance on the AHA dataset [11], which includes (1) unseen simulated environments and (2) unseen error category (control errors), even surpassing the VLM-based baseline trained directly on these tasks by +19 points in accuracy. Furthermore, I-FailSense transfers to real-world semantic misalignment failure detection with minimal fine-tuning of the FS blocks, achieving 74% accuracy. These results demonstrate the importance of designing architectures for robotic failure detection that effectively leverage the internal representations learned by VLMs. They also highlight that semantic misalignment failures are a particularly challenging and informative class of language-conditioned robotic manipulation errors.
@@ IV Method
In this work, we focus on detecting language-conditioned robotic manipulation failures, with an emphasis on semantic misalignment errors. We introduce datasets targeting this challenging failure detection problem (Section IV-A), and propose a new method for post-training a VLM agent and leveraging its internal representation for failure detection (Section IV-B). The overall architecture is presented in Figure 2.
@@ IV-B1 Supervised Fine-Tuning (Stage 1)
$\mathcal{L}_{\text{CE}}(\theta)=-\frac{1}{M}\sum_{i=1}^{M}\log p_{\theta}\left(t_{i}\mid x,t_{<i}\right),$ (1)
where $M$ is the sequence length, $t_{i}$ is the target token at position $i$, $t_{<i}$ represents all previous tokens, $x$ is the input context (i.e. the prompt), and $p_{\theta}(\cdot)$ denotes the probability distribution of the model with parameters $\theta$, over the vocabulary.
@@ V Experimental Setup
We train I-FailSense on a semantic misalignment dataset, built following the pipeline presented in Section IV-B. We design experiments addressing the following research questions:
  • Q1
    Can I-FailSense reliably detect semantic misalignment errors?
  • Q2
    Does this capability transfer to detecting control errors?
  • Q3
    How well does I-FailSense generalize to failure detection on out-of-distribution (OOD) simulated trajectories?
  • Q4
    How well does it generalize to failure detection on real-world ?
@@ VI-D Generalization to Real-World Trajectories (Q4)
TABLE II: Evaluation metrics of I-FailSense on $\mathcal{D}_{\text{SMF-DROID}}$ with different fine-tuning of the FS blocks on $\mathcal{D}_{\text{SMF-CALVIN}}$ and/or $\mathcal{D}_{\text{SMF-DROID}}$ compared with zero-shot VLMs. The FS blocks training dataset is put into brackets. Best results are shown in bold.
Table II shows the performance of I-FailSense evaluated on the $\mathcal{D}_{\text{SMF-DROID}}$ dataset, compared to zero-shot VLMs of different sizes. The I-FailSense models are trained with LoRA on $\mathcal{D}_{\text{SMF-CALVIN}}$, while the FS blocks are trained on $\mathcal{D}_{\text{SMF-CALVIN}}$ and/or $\mathcal{D}_{\text{SMF-DROID}}$. The results demonstrate that directly transferring I-FailSense fully trained on simulated trajectories ($\mathcal{D}_{\text{SMF-CALVIN}}$) to detect failures in real-world trajectories only partially succeeds: it improves the accuracy but decreases F1 score due to a significant drop in recall. This indicates that a model trained solely on simulated data struggles to recognize failures in real-world trajectories, even though precision increases, meaning that it better identifies successful trajectories. Training the FS blocks on both simulated and real-world trajectories ($\mathcal{D}_{\text{SMF-CALVIN}}$ + $\mathcal{D}_{\text{SMF-DROID}}$) reduces the recall drop (from -0.4 to -0.23) while increasing precision ( +0.27), surpassing the base model in both accuracy and F1 score. The best performance is achieved by training the FS blocks solely on real-world failure detection ($\mathcal{D}_{\text{SMF-DROID}}$), suggesting that training on simulated data can hinder learning patterns relevant for real-world errors detection. Nevertheless, I-FailSense still benefits from the LoRA representation learned during the initial training, enabling effective failure detection in real-world trajectories. In fact, the top-performing I-FailSense matches the performance of the best zero-shot VLM, Qwen2.5-VL, despite having half the number of parameters. Interestingly, while I-FailSense performs better using only the exocentric PoV compared to the dual-PoV model (+0.02 points in accuracy) for failure detection in simulated trajectories, on real-world trajectories the combination of exocentric and egocentric PoV achieves the best performance, surpassing the single-PoV model by +0.03 points in accuracy. This can be explained by the higher number of distractors and less stable viewpoints in real-world trajectories as depicted in Figure 5, compared to the more controlled simulation environments (Figures 3 and 4). These factors make failure detection more challenging in real-world settings, hence the need for two PoV.
@@ Acknowledgments
Experiments presented in this paper were carried out using the HPC resources of IDRIS under the allocation 2025-[AD011015740] made by GENCI. This work was supported by the European Commission’s Horizon Europe Framework Programme under grant No 101070381 (PILLAR-robots). This research was partially funded by the French National Research Agency (ANR) under the OSTENSIVE projects (ANR-24-CE33-6907-01), ANITA project (ANR-22-CE38-0012-01) and the France 2030 program, reference ANR-23-PAVH-0005 (INNOVCARE Project).

### v3

@@ I Introduction
Empirically, we demonstrate that I-FailSense outperforms zero-shot SOTA VLMs in detecting semantic misalignment errors reaching an accuracy of 90% in our proposed datasets built from simulated trajectories. Remarkably, I-FailSense generalizes beyond its training data and achieves strong performance on the AHA dataset [5], which includes (1) unseen simulated environments and (2) unseen error category (control errors), even surpassing the VLM-based baseline trained directly on these tasks by +19 points in accuracy. Furthermore, I-FailSense transfers to real-world semantic misalignment failure detection with minimal fine-tuning of the FS blocks, achieving 74% accuracy.
@@ IV Method
In this work, we focus on detecting language-conditioned robotic manipulation failures, with an emphasis on semantic misalignment errors. We introduce datasets targeting this challenging failure detection problem (Section IV-A), and propose a new method for post-training a VLM agent and leveraging its internal representation for failure detection (Section IV-B). Figure 2 presents the overall architecture.
@@ IV-B1 Supervised Fine-Tuning (Stage 1)
$\mathcal{L}_{\text{CE}}(\theta)=-\log p_{\theta}\left(t^{*}\mid x\right),$ (1)
where $t^{*}$ is the target token, $x$ is the input context (i.e. the prompt), and $p_{\theta}(\cdot)$ denotes the probability distribution of the model with parameters $\theta$, over the vocabulary.
@@ V Experimental Setup
We train I-FailSense on a semantic misalignment dataset, built following the pipeline presented in Section IV-B. We design experiments addressing the following research questions: Q1: Can I-FailSense reliably detect semantic misalignment errors? Q2: Does this capability transfer to detecting control errors? Q3: How well does I-FailSense generalize to failure detection on out-of-distribution (OOD) simulated trajectories? Q4: How well does it generalize to failure detection on real-world ?
@@ VI-D Generalization to Real-World Trajectories (Q4)
TABLE II: Evaluation metrics of I-FailSense on $\mathcal{D}_{\text{SMF-DROID}}$ with different fine-tuning of the FS blocks on $\mathcal{D}_{\text{SMF-CALVIN}}$ and/or $\mathcal{D}_{\text{SMF-DROID}}$ compared with zero-shot VLMs. FS blocks training dataset is into brackets, best results in bold.
@@
Table II shows the performance of I-FailSense evaluated on the $\mathcal{D}_{\text{SMF-DROID}}$ dataset, compared to zero-shot VLMs of different sizes. The I-FailSense models are trained with LoRA on $\mathcal{D}_{\text{SMF-CALVIN}}$, while the FS blocks are trained on $\mathcal{D}_{\text{SMF-CALVIN}}$ and/or $\mathcal{D}_{\text{SMF-DROID}}$. The results demonstrate that directly transferring I-FailSense fully trained on simulated trajectories ($\mathcal{D}_{\text{SMF-CALVIN}}$) to detect failures in real-world trajectories only partially succeeds: it improves the accuracy but decreases F1 score due to a significant drop in recall. This indicates that a model trained solely on simulated data struggles to recognize failures in real-world trajectories, even though precision increases, meaning that it better identifies successful trajectories. Training the FS blocks on both simulated and real-world trajectories ($\mathcal{D}_{\text{SMF-CALVIN}}$ + $\mathcal{D}_{\text{SMF-DROID}}$) reduces the recall drop (from -0.4 to -0.23) while increasing precision ( +0.27), surpassing the base model in both accuracy and F1 score. The best performance is achieved by training the FS blocks solely on real-world failure detection ($\mathcal{D}_{\text{SMF-DROID}}$), suggesting that training on simulated data can hinder learning patterns relevant for real-world errors detection. Nevertheless, I-FailSense still benefits from the LoRA representation learned during the initial training, enabling effective failure detection in real-world trajectories. In fact, the top-performing I-FailSense matches the performance of the best zero-shot VLM, Qwen2.5-VL, despite having half the number of parameters. Interestingly, while I-FailSense performs better using only the exocentric PoV compared to the dual-PoV model (+0.02 points in accuracy) for failure detection in simulated trajectories, on real-world trajectories the combination of exocentric and egocentric PoV achieves the best performance, surpassing the single-PoV model by +0.03 points in accuracy. This can be explained by the higher number of distractors and less stable viewpoints in real-world trajectories as depicted in Figure 5, compared to the more controlled simulation environments (Figures 3 4). These factors make failure detection more challenging in real-world settings, hence the need for two PoV.
@@ Acknowledgments
This work used IDRIS HPC resources under the allocation 2025-[AD011015740] made by GENCI, was supported by the European Commission’s Horizon Europe Framework Programme under grant No 101070381 (PILLAR-robots), and was partially funded by the French National Research Agency (ANR) under the OSTENSIVE projects (ANR-24-CE33-6907-01), ANITA project (ANR-22-CE38-0012-01) and the France 2030 program, reference ANR-23-PAVH-0005 (INNOVCARE Project).